OpenStudy (anonymous):

Use a sum or difference identity to find an exact value sin(15 degrees)

OpenStudy (anonymous):

15 is half of 30 so use half angle formula for sine

OpenStudy (zehanz):

This will be a rather long story... Note: everything becomes a lot clearer if you write this out yourself, with proper rationals, and not everything on one and the same line... 30 + 15 = 45. We know the exact values of sin and cos for 30 and 45 degrees. Also the formulas for sum of angles are sin 45 = sin(30+15) = sin30cos15 + cos30sin15. (i) cos 45 = cos(30+15) = cos30cos15 - sin30sin15. (ii) Now, sin45 = cos45 (= sqrt(2)/2), so (i) = (ii). Also, sin30 = 1/2 and cos30 = sqrt(3)/2. Putting these values in (i) and (ii) gives: 1/2 cos15 + sqrt(3)/2 sin15 = sqrt(3)/2 cos15 - 1/2 sin15. Collect terms with sin15 and with cos15: (sqrt(3)/2 + 1/2)sin15 = (sqrt(3)/2 - 1/2)cos15. Divide by sqrt(3)/2 + 1/2: sin15 = (sqrt(3)/2 - 1/2)/(sqrt(3)/2 + 1/2) * cos15. Isolate the factor 1/2: sin15 = (1/2 *(sqrt(3) - 1))/(1/2 * (sqrt(3) + 1) * cos15. Now get rid of the 1/2: sin15 = (sqrt(3) - 1)/(sqrt(3) +1) * cos15. (iii) We seem to be stuck at this point, because now we also need cos15 :(. Now remember the formula sin2x=2sinxcosx. This means that sin30=2sin15*cos15. So cos15 = sin30/(2sin15) = 1/2/(2sin15) = 1/(4sin15). We can use this result in (iii) and get: sin15 = (sqrt(3) - 1)/(sqrt(3) + 1) * (1/(4sin15)). Multiply both sides with 4sin15: 4(sin15)^2 = (sqrt(3) - 1)/(sqrt(3) + 1). (iv) Now it is time to "simplify" the right side of (iv), by multiplying both nominator and denominator with sqrt(3) - 1. 4(sin15)^2 = (sqrt(3) - 1)^2 / ((sqrt(3) + 1)(sqrt(3) - 1)) 4(sin15)^2 = (3 - 2sqrt(3) +1) / (3 - 1) = (4 - 2sqrt(3)) / 2 = 2 - sqrt(3). (v) Hold on, we are almost there! In (v), divide by 4 and take the root (sin15 is in the first quadrant, so it is positive and we do not need to worry about a negative value...). This gives: sin15 = sqrt((2 - sqrt(3))/4). Taking the root of 4 gives, at last, sin15 = sqrt(2 - sqrt(3)) / 2 Hope this helps! ZeHanz

OpenStudy (unklerhaukus):

\[\sin(15°)\]\[=\sin(45°-30°)\] \[\qquad\qquad\qquad \boxed{\sin(u\pm v)=\sin(u)\cos(v)\pm \cos(u)\sin(v)}\] \[=\sin(45°)\cos(30°)-\cos(45°)\sin(30°)\]   \[=\frac{\sqrt 2}2\cdot\frac{\sqrt3}2-\frac{\sqrt 2}2\cdot\frac12\]\[=\frac{\sqrt2(\sqrt 3-1)}4\]